A lighthouse keeper observes that there is a 3 degree angle of depression between the horizontal and the line of sight to a ship. If the keeper is 19 m above the water how far is the ship from shore?
this unit is called solving right triangles.
please help if you can. the angle of depression part is confusing me. can you explain?

Answer by scott8148(6628)

the angle of depression means that the keeper looks 3 degrees down from the horizontal to see the ship
__ this means that someone on the ship looks 3 degrees up from the horizontal to see the keeper

the horizontal and the lighthouse form a right angle
__ the distance from the keeper to the ship is the hypotenuse
__ the lighthouse is the side opposite the 3° angle
__ the distance from the ship to the shore is the side adjacent to the 3° angle

let x="distance to shore"

19m/x=tan(3°) __ 19m/(tan(3°))=x